a metal strip 6.5cm long,0.88cm wide &0.76mm thick moves with constant velocity v through a magnetic field B=1.2mT perpendicular to the strip.A Potential difference of 3.9V is measured between points x & y across the strip.Calculate the speed V.
1
Expert's answer
2019-05-03T09:06:26-0400
"v= \frac { E_H} { B} = = \frac { \frac { ΔV_H} { w}} { B} (1)"

where ΔVH=3.9V, w=0.88×10-2 m, B=1.2×10-3 T

Using (1) we got:

v=3.69×105 m/s